OpenStudy (anonymous):

-4y + 17 29 Part 1: Solve the inequality above. Part 2: Describe the graph of the solution.

OpenStudy (anonymous):

can you write the inequality sign in the equation plz :)

OpenStudy (anonymous):

5.01] -4y + 17=< 29 Part 1: Solve the inequality above. Part 2: Describe the graph of the solution.

OpenStudy (jdoe0001):

what would you get if you were to solve for "y" on -> -4y + 17= 29 ?

OpenStudy (anonymous):

i dont know

OpenStudy (jdoe0001):

so, do you know how to factor? because if you don't, I don't see how any of this would apply

OpenStudy (anonymous):

?

OpenStudy (anonymous):

factor ?

OpenStudy (jdoe0001):

well, why give you an exercise of something you haven't covered yet

OpenStudy (anonymous):

@joshuapangilinan123 what do you get if you subtract 17 from both side of -4y + 17 =29